Timberly Construction makes a lump-sum purchase of several assets on January 1 at a total cash price of $810,000.The estimated market values of the purchased assets are building, $427,800; land, $241,,800; land improvements, $37,200; and four vehicles, $223,200

Compute the first year depreciation expense on the land improvements assuming a five year and double declining balance depreciation

Depreciation expense on land improvements?
Answer #1

Answer

  • Total estimated market values = 427800 + 241800 + 37200+ 223200= $ 930,000
  • Market value of Land Improvement = $ 37,200
  • Cost allocated to Land Improvement = $ 810000 lumpsum paid x $ 37200/ 930000
    = 810000 x 37200 / 930000
    = $ 32,400
  • Answer calculation:

A

Cost

$            32,400.00

B

Residual Value

$                           -  

C=A - B

Depreciable base

$            32,400.00

D

Life [in years]

5

E=C/D

Annual SLM depreciation

$              6,480.00

F=E/C

SLM Rate

20.00%

G=F x 2

DDB Rate

40.00%

>First Year Depreciation Expense = $ 32,400 x 40% = $ 12,960 Answer